OpenStudy (anonymous):

suppose you have a bag that holds 30 pieces of candy:11 blue berry, 5lemon and 14 cream soda you reach into the bag to pull out one piece find the probability that the piece you pull out is not cream soda flavored.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how many pieces are not cream flavor?

jimthompson5910 (jim_thompson5910):

there's only 30 pieces total

jimthompson5910 (jim_thompson5910):

any idea?

OpenStudy (anonymous):

16 14/16? sorry misread ive been doing math for 2hours straight.

jimthompson5910 (jim_thompson5910):

16 non cream flavor out of 30 total \[\Large P(\text{not cream flavored})=\frac{\text{number of non-cream flavor}}{\text{number total}}\] \[\Large P(\text{not cream flavored})=\frac{16}{30}\] \[\Large P(\text{not cream flavored})=\frac{8}{15}\]

jimthompson5910 (jim_thompson5910):

Or you can do it this way \[\Large P(\text{not cream flavored})=1 - P(\text{Cream flavor})\] \[\Large P(\text{not cream flavored})=1 - \frac{\text{number of cream flavor}}{\text{number total}}\] \[\Large P(\text{not cream flavored})=1 - \frac{14}{30}\] \[\Large P(\text{not cream flavored})=\frac{30}{30} - \frac{14}{30}\] \[\Large P(\text{not cream flavored})=\frac{30-14}{30}\] \[\Large P(\text{not cream flavored})=\frac{16}{30}\] \[\Large P(\text{not cream flavored})=\frac{8}{15}\] So you get the same answer either way

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

no problem